# Sortability of Time Series Data

### v2

@@ Abstract
Evaluating the performance of causal discovery algorithms that aim to find causal relationships between time-dependent processes remains a challenging topic. In this paper, we show that certain characteristics of datasets, such as varsortability Reisach et al. [2021] and $R^{2}$-sortability Reisach et al. [2023], also occur in datasets for autocorrelated stationary time series. We illustrate this empirically using four types of data: simulated data based on SVAR models and Erdős-Rényi graphs, the data used in the 2019 causality-for-climate challenge [Runge et al., 2019], real-world river stream datasets, and real-world data generated by the Causal Chamber of [Gamella et al., 2024]. To do this, we adapt var- and $R^{2}$-sortability to time series data. We also investigate the extent to which the performance of score-based causal discovery methods goes hand in hand with high sortability. Arguably, our most surprising finding is that the investigated real-world datasets exhibit high varsortability and low $R^{2}$-sortability indicating that scales may carry a significant amount of causal information.
@@ 1 Introduction
Inferring causal relationships between variables in a multivariate time series setting is an ongoing topic of research in many fields, including economics/econometrics [Varian, 2016], climate science [Runge et al., 2019, Runge, 2020], medicine [Yazdani and Boerwinkle, 2015] and neuroscience [Bergmann and Hartwigsen, 2021]. The process of inferring causal structure from data is often referred to as causal discovery or causal structure learning. Most of the literature on causal discovery is devoted to the following two types of methods: Constraint-based causal structure learning algorithms, such as the PC-algorithm [Spirtes and Glymour, 1991] and PCMCI [Runge, 2020] for time series data, use tests of conditional independence to iteratively learn a causal graph. Score-based method such as GES [Chickering, 2002] fit a directed acyclic graph (DAG) to the data by optimising a score function, but need to search a large discrete space of DAGs which is an NP-hard problem.
@@
More recently, [Zheng et al., 2018] have proposed to embed the discrete DAG-search space into a continuous one through a differentiable acyclicity constraint to make the problem amenable to continuous (gradient based) optimisation. The method introduced in [Zheng et al., 2018], NOTEARS, showed impressive performance on data simulated with linear additive noise models (LANMs). However, as shown by Reisach et al. [2021] the strong performance of NOTEARS and similar methods on LANM data vanished after the data was normalized. Reisach et al. [2021] noticed that, before normalization, additive noise model data is highly varsortable, meaning that, on average, the causal order of the system can be recovered well by sorting the variables by the amplitude of their estimated variances. Since high varsortability and good NOTEARS performance were highly correlated, they therefore conjectured that NOTEARS implicitly made use of variance sorting in its optimisation. The issue has recently been revisited by Ng et al. [2024], who pointed out that (a) NOTEARS does not necessarily perform well in the presence of high varsortability and that (b) normalisation of data generated by LANMs moves the data far away from the assumption of equal noise variances that underlies the NOTEARS methodology. Thus, according to Ng et al. [2024], NOTEARS should not have been expected to perform well on normalized LANM data in the first place as one of its fundamental assumptions is not satisfied.
@@
Rather than weighing in on this discussion, in this workshop contribution we explore how much var- and $R^{2}$-sortability (another sortability criterion introduced in [Reisach et al., 2023]) can be seen in data commonly used in method validation of time series causal discovery methods. We evaluate the degree of var-/$R^{2}$-sortability in different datasets as well as the performance of different algorithms for time series causal discovery in the presence/absence of var-/$R^{2}$-sortability and before/after normalisation. One of the main questions underlying the discussions on sortability is this: how much var-/$R^{2}$-sortability do we expect to see in real-world data? The answer to this question is likely highly context-dependent, and the question can be notoriously difficult to settle even for a single dataset, since one needs to know the causal ground truth to compute sortability. We investigate sortability scores in two of the rare cases where a causal ground-truth is available, the river flow dataset used in Tran et al. [2024] and the Causal Chamber datasets recently published by Gamella et al. [2024]. In both cases, we find sortability values quite far from $0.5$. In the river stream datasets, marginal variances tend to decrease along the causal order (varsortability close to zero) while in the Causal Chamber case, marginal variances increase along the causal order (varsortability close to one). For the $R^{2}$-score we find the exact opposite: values close to one for some rivers in the river dataset, and values close to zero for some of the causal chamber datasets.
@@
These observations illustrate that discarding scales in causal discovery as arbitrary may be premature, as scales may encode significant causal information. We hypothesise plausible physical explanations for the observed varsortability in the investigated datasets which also call into question the validity of an equal noise variance assumption in these cases.
@@
    We extend var- and $R^{2}$-sortability and the simple benchmarking algorithm of Reisach et al. [2021, 2023] to the time series setting.
@@
    We demonstrate that there is a positive correlation between the performance of score-based causal discovery algorithms for time series and the varsortability of data generated by structural autoregressive processes.
@@
    We investigate sortability of the data used in the 2019 causal discovery challenge [^1] https://causeme.uv.es/neurips2019/ by Runge et al. [2019].
@@
    We calculate var- and $R^{2}$-sortability of the river flow dataset used in Tran et al. [2024], and of the recently published datasets generated by the Causal Chamber[Gamella et al., 2024]. We find that in these real-live datasets, the scale plays an important role for potential causal discovery algorithms.
@@ 2.1 Causal discovery for time series
A stationary time series graph (ts-graph) is a directed graph $\mathcal{G}=(V\times\mathbb{Z},\mathcal{D}),\ V=\{1.\dots,d\}$, whose edges $((i,t-k),(j,t))$ are assumed invariant under translation of the time component. In addition, it is typically required that there is a finite maximal lag $\tau_{max}=\max_{i,j\in V}\{k|((i,t-k),(j,t))\in\mathcal{D}\}<\infty$ and that the contemporaneous component of $\mathcal{G}$ is acyclic. Any stationary ts-graph induces a directed, potentially cyclic summary graph $\mathcal{G}_{\mathrm{sum}}$ over $V$ that contains a directed edge $(i,j)$ if $(i,t-k)\to(j,t)\in\mathcal{D}$ for some $k$. The adjacency matrices of $\mathcal{G},\mathcal{G}_{\mathrm{sum}}$ will be denoted by $\mathbf{W},\mathbf{W}_{\mathrm{sum}}$ respectively.
The aim of most causal discovery methods for time series is to recover either $\mathcal{G}$ or $\mathcal{G}_{\mathrm{sum}}$ from observational or interventional data. Often the data is assumed to be generated by a discrete multivariate process $(\mathbf{X}_{t})_{t\in\mathbb{Z}},\ \mathbf{X}_{t}=(X_{t}^{1},...,X_{t}^{(d)})$ compatible with $\mathcal{G}$. The process $(\mathbf{X}_{t})_{t\in\mathbb{Z}}$ is typically modelled as a structural vector autoregressive (SVAR) process, in which case, compatibility with $\mathcal{G}$ means that $(\mathbf{X}_{t})_{t\in\mathbb{Z}}$ follows the evolution rule
@@
for $j\in\{1,...,d\}$, where $\eta^{j}$ are Gaussian white noise processes, $\mathrm{pa}_{\mathcal{G}_{\mathrm{sum}}}(j)$ denotes the parents of node $j$ in the summary graph, and $a^{j}_{i,t-k}$ is only allowed to be non-zero if $(i,t-k)\to(j,t)$ is an edge in $\mathcal{G}$. SVAR-processes can be considered the time series analogue of additive linear noise models for which sortability was discussed in [Reisach et al., 2021, 2023]. When generating data from such a model, coefficients are typically randomly drawn, sometimes with a pre-specified proportion of contemporaneous links, and the process is being run until it has converged to a stationary distribution (or is discarded if the distribution is non-stationary).
@@ 2.2 NOTEARS and Derivatives
Zheng et al. [2018] propose the score-based causal discovery method NOTEARS, which embeds the discrete search space of DAGs into a continuous one by using the differentiable function $h(\mathbf{W})=tr\>e^{\mathbf{W}\circ\mathbf{W}}-d$. This function is 0 if and only if $\mathbf{W}$ is the adjacency matrix of an acyclic graph and hence measures the DAGness of $\mathbf{W}$ [Zheng et al., 2018]. By combining this function with a score evaluating how well the estimated weight matrix $\mathbf{W}$ fits the data, Zheng et al. [2018] formulate the constrained optimisation problem to find
$\min_{\mathbf{W}}\frac{1}{n}||\mathbf{X}-\mathbf{X}\mathbf{W}||^{2}_{2}$ (2)
@@
The DYNOTEARS algorithm Pamfil et al. [2020] modifies the NOTEARS algorithm to work with time-lagged and auto-correlated dependencies by redefining the optimisation problem to $\min_{\mathbf{W}1}\ell(\mathbf{W_{c}},\mathbf{W_{l}})\text{ s.t. }\mathbf{W}_{c}\text{ is acyclic}$, where $\mathbf{W_{l}}\in\mathbb{R}^{d\times d\times\tau_{max}}$ is the lagged adjacency matrix and $\mathbf{W}_{c}$ is the contemporaneous adjacency of the underlying time series process. As the underlying time series graph to estimate is acyclic if and only if $\mathbf{W}_{c}$ is acyclic, it suffices to enforce the acyclicity constraint only on $\mathbf{W}_{c}$ [Pamfil et al., 2020]. To ensure sparsity of $\mathbf{W}$, Pamfil et al. [2020] also add an $\ell_{1}$ penalty term, leading to the constraint optimisation problem
$\min_{\mathbf{W}}\frac{1}{2n}||\mathbf{X}-\mathbf{X}\mathbf{W}_{c}-\mathbf{X_{l}}\mathbf{W}_{l}||^{2}_{2}+\lambda_{1}||\mathbf{W}_{c}||_{1}+\lambda_{2}||\mathbf{W}_{l}||_{1}$ (3)
@@
The continuous optimisation problems of NOTEARS and DYNOTEARS can be solved efficiently by rewriting the problem using the augmented Lagrangian method and using a numerical solver such as L-BFGS [Zheng et al., 2018, Pamfil et al., 2020]. After applying the numerical optimisation algorithm in both of the algorithms, a threshold $t$ is applied to remove weights close to zero [Zheng et al., 2018, Pamfil et al., 2020].
@@ 2.3 Sortability Criteria
In the following, we briefly reiterate the two sorting criteria varsortability and $R^{2}$-sortability introduced by Reisach et al. [2021] and Reisach et al. [2023]. In essence, both of these approaches calculate a score $s$ of sortability in a comparable manner; $s$ is the measurement of the degree of agreement between the true causal order and the increasing order of a sortability criterion $cri$.
For any causal model containing the variables $\{X^{(1)},\ldots,X^{(d)}\}$ with a (non-degenerate) adjacency matrix $\mathbf{W}$, the sortability score is the fraction of directed paths that start from a node with a strictly lower sortability criterion than the node they end in. Thus the sortability for one selected criterion $cri$ can be calculated as
@@
In the case of varsortability Reisach et al. [2021], the criterion $cri(X^{i})=\mathrm{Var}(X^{i})$ is the marginal variance, whereas in the case of $R^{2}$-sortability, it is the coefficient of determination $cri(X^{i})=R^{2}(X^{i})$ which acts as a proxy for the fraction of the variance of $X_{i}$ that is explained by its causal parents, see [Reisach et al., 2023] for details. Varsortalility $v$ is defined by using the marginal variance as $cri$. $R^{2}$-sortability $r$ is defined by using the obtained $R^{2}$-coefficients as the sorting criterion $cri$ [Reisach et al., 2021, 2023].
Reisach et al. [2021] showed that varsortability is usually high in data generated by LANMs (see the analytical and empirical proof in their paper). Based on their sortability criteria, Reisach et al. [2021] and Reisach et al. [2023] introduced the baseline methods varsortnregress and $R^{2}$-sortnregress respectively. These algorithms sort the system variables based on their variances or $R^{2}$-scores which are estimated by fitting a regression model; these simple benchmark methods are then shown to have similar performance to some state of the art causal discovery algorithms [Reisach et al., 2021, 2023] on LANM data.
@@ 3 Modified Sortability Criteria for Summary Graphs
However, any pair of processes $X^{i}$ and $X^{j}$ that belong to the same strongly connected component of the summary graph (i.e. that are connected by a cycle) would always contribute a $1$ to the denominator and a $2$ to the numerator of Equation 4.
@@
$s:=\frac{\sum_{(i,j)\in\mathcal{AP}(\mathcal{G}_{\mathrm{sum}})}increasing(cri(X_{i}),cri(X_{j}))}{\sum^{d-1}_{k=1}\sum_{(i,j)\in\mathcal{AP}(\mathcal{G}_{\mathrm{sum}})}1}\in[0,1],$ (6)
@@
$C$$B$$D$$A$$Cri(C)=3$$Cri(B)=2$$Cri(D)=2.5$$Cri(A)=1$
Figure 1: Example of the calculation of $s$ with cycles. The contribution of the pairs $(B,C)$ and $(C,B)$ is ignored since the two nodes belong to the black cycle. Cycle-free directed paths that connect admissible node pairs are depicted in colour.
For the example in Figure 1, the sortability score $s$ is $\frac{{\color[rgb]{1,0.6015625,0.6015625}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.6015625,0.6015625}1}+{\color[rgb]{1,0.703125,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.703125,0.3984375}1}+{\color[rgb]{0.3125,0.66015625,0.328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.3125,0.66015625,0.328125}1}+{\color[rgb]{0.3984375,0.3984375,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.3984375,0.3984375,1}0}}{{\color[rgb]{1,0.6015625,0.6015625}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.6015625,0.6015625}1}+{\color[rgb]{1,0.703125,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.703125,0.3984375}1}+{\color[rgb]{0.3125,0.66015625,0.328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.3125,0.66015625,0.328125}1}+{\color[rgb]{0.3984375,0.3984375,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.3984375,0.3984375,1}1}}=0.75$. If we would calculate it including cyclicly connected pairs, it would result in $\frac{4}{6}\approx 0.67$.
@@
Based on heuristics and empirical findings, Reisach et al. [2021] argue that high variability causes gradient-based optimisation algorithms such as NOTEARS[Zheng et al., 2018] to favour graphs whose edges point in the direction of increasing marginal variances during their first optimisation step. Since DYNOTEARS is a time-series adaptation of NOTEARS, that essentially coincides with NOTEARS on the contemporaneous components of a ts-graph, [Reisach et al., 2021]’s arguments might be applicable to (the contemporaneous part of) DYNOTEARS as well.
However, in the recent work Ng et al. [2024] provide examples in which continuous optimization-based methods can not perform well in the presence of high varsortability. Ng et al. [2024] also give an alternative explanation for why continuous structure learning performs significantly worse after standardisation of LANM data. Continuous structure learning assumes equal noise variances for all variables in the system, which is violated in the standardised model leading to poor performance.

### v3

@@ Abstract
Evaluating the performance of causal discovery algorithms that aim to find causal relationships between time-dependent processes remains a challenging topic. In this paper, we show that certain characteristics of datasets, such as varsortability Reisach et al. (2021) and $R^{2}$-sortability Reisach et al. (2023), also occur in datasets for autocorrelated stationary time series. We illustrate this empirically using four types of data: simulated data based on SVAR models and Erdős-Rényi graphs, the data used in the 2019 causality-for-climate challenge (Runge et al., 2019), real-world river stream datasets, and real-world data generated by the Causal Chamber of Gamella et al. (2024). To do this, we adapt var- and $R^{2}$-sortability to time series data. We also investigate the extent to which the performance of continuous score-based causal discovery methods goes hand in hand with high sortability. Arguably, our most surprising finding is that the investigated real-world datasets exhibit high varsortability and low $R^{2}$-sortability indicating that scales may carry a significant amount of causal information.
@@ 1 Introduction
Inferring causal relationships between variables in a multivariate time series setting is an ongoing topic of research in many fields, including economics/econometrics (Varian, 2016), climate science (Runge et al., 2019; Runge, 2020), medicine (Yazdani & Boerwinkle, 2015) and neuroscience (Bergmann & Hartwigsen, 2021). The process of inferring causal structure from data is often referred to as causal discovery or causal structure learning. Most of the literature on causal discovery is devoted to the following two types of methods: Constraint-based causal structure learning algorithms, such as the PC-algorithm (Spirtes & Glymour, 1991) and PCMCI (Runge, 2020) for time series data, use tests of conditional independence to iteratively learn a causal graph. Score-based method such as GES (Chickering, 2002) fit a directed acyclic graph (DAG) to the data by optimising a score function, but need to search a large discrete space of DAGs which is an NP-hard problem.
@@
More recently, Zheng et al. (2018) have proposed to embed the discrete DAG-search space into a continuous one through a differentiable acyclicity constraint to make the problem amenable to continuous (gradient based) optimisation. The method introduced in Zheng et al. (2018), NOTEARS, showed impressive performance on data simulated with linear additive noise models (LANMs). However, as shown by Reisach et al. (2021) the strong performance of NOTEARS and similar methods on LANM data vanished after the data was normalized. Reisach et al. (2021) noticed that, before normalization, additive noise model data is highly varsortable, meaning that, on average, the causal order of the system can be recovered well by sorting the variables by the amplitude of their estimated variances. Since high varsortability and good NOTEARS performance were highly correlated, they therefore conjectured that NOTEARS implicitly made use of variance sorting in its optimisation. The issue has recently been revisited by Ng et al. (2024), who pointed out that (a) NOTEARS does not necessarily perform well in the presence of high varsortability and that (b) normalisation of data generated by LANMs moves the data far away from the assumption of equal noise variances that underlies the NOTEARS methodology. Thus, according to Ng et al. (2024), NOTEARS should not have been expected to perform well on normalized LANM data in the first place as one of its fundamental assumptions is not satisfied.
@@
Rather than weighing in on this discussion, in this contribution we explore how much var- and $R^{2}$-sortability (the coefficient of determination which acts as a proxy for the fraction of the variance a variable that is explained by its causal parents introduced in Reisach et al. (2023)) can be seen in data commonly used in method validation of time series causal discovery methods. We evaluate the degree of var-$R^{2}$-sortability in different datasets as well as the performance of different algorithms for time series causal discovery in the presence/absence of var-$R^{2}$-sortability and before/after normalisation. One of the main questions underlying the discussions on sortability is this: how much var-$R^{2}$-sortability do we expect to see in real-world data? The answer to this question is likely highly context-dependent, and the question can be notoriously difficult to settle even for a single dataset, since one needs to know the causal ground truth to compute sortability. We investigate sortability scores in two of the rare cases where a causal ground-truth is available, the river flow dataset used in Tran et al. (2024) and the Causal Chamber datasets recently published by Gamella et al. (2024). In both cases, we find sortability values quite far from $0.5$ (the case where there is no information in the sortability criterion), meaning that there is information in the variance or $R^{2}$-values of the data. In the river stream datasets, marginal variances tend to decrease along the causal order (varsortability close to zero) while in the Causal Chamber case, marginal variances increase along the causal order (varsortability close to one). For the $R^{2}$-score we find the exact opposite: values close to one for some rivers in the river dataset, and values close to zero for some of the causal chamber datasets.
@@
These observations illustrate that discarding scales in causal discovery as arbitrary may be premature, as scales may encode significant causal information. We hypothesise plausible physical explanations for the observed varsortability in the investigated datasets which also call into question the validity of an equal noise variance assumption in these cases. For the river data we know that the the width of the rivers decrease from the source to the mouth which has potentially an influence on the variance of extreme flows.
@@
    We extend var- and $R^{2}$-sortability and the simple benchmarking algorithm of Reisach et al. (2021; 2023) to the time series setting.
@@
    We demonstrate that there is a positive correlation between the performance of continuous score-based causal discovery algorithms for time series and the varsortability of data generated by structural autoregressive processes.
@@
    We investigate sortability of the data used in the 2019 causal discovery challenge [^1] https://causeme.uv.es/neurips2019/ by Runge et al. (2019) and show that some data sets are highly varsortable and simple benchmark perform well on these.
@@
    We calculate var- and $R^{2}$-sortability of the river flow dataset used in Tran et al. (2024), and of the recently published datasets generated by the Causal Chamber(Gamella et al., 2024). We find that in these real-live datasets, the scale plays an important role for potential causal discovery algorithms.
@@ 2.1 Causal discovery for time series
A stationary time series graph (ts-graph) is a directed graph $\mathcal{G}=(V\times\mathbb{Z},\mathcal{D}),\ V=\{1.\dots,d\}$, whose edges $(i,t-k)\to(j,t)$ are assumed invariant under translation of the time component. In addition, it is typically required that there is a finite maximal lag $\tau_{max}=\max_{i,j\in V}\{k|((i,t-k),(j,t))\in\mathcal{D}\}<\infty$ and that the contemporaneous component of $\mathcal{G}$ is acyclic, where $k\in\mathbb{N}$. Any stationary ts-graph induces a directed, potentially cyclic summary graph $\mathcal{G}_{\mathrm{sum}}$ over $V$ that contains a directed edge $(i,j)$ if $(i,t-k)\to(j,t)\in\mathcal{D}$. The adjacency matrices of $\mathcal{G},\mathcal{G}_{\mathrm{sum}}$ will be denoted by $\mathbf{W},\mathbf{W}_{\mathrm{sum}}$ respectively.
The aim of most causal discovery methods for time series is to recover either $\mathcal{G}$ or $\mathcal{G}_{\mathrm{sum}}$ from observational or interventional data. Often the data is assumed to be generated by a discrete multivariate process $(\mathbf{X}_{t})_{t\in\mathbb{Z}},\ \mathbf{X}_{t}=(X_{t}^{1},...,X_{t}^{(d)})$ compatible with $\mathcal{G}$. The process $(\mathbf{X}_{t})_{t\in\mathbb{Z}}$ is typically modelled as a structural vector autoregressive (SVAR) process(Hyvärinen et al., 2010), in which case, compatibility with $\mathcal{G}$ means that $(\mathbf{X}_{t})_{t\in\mathbb{Z}}$ follows the evolution rule
@@
for $j\in\{1,...,d\}$, where $\eta^{j}$ are Gaussian white noise processes, $\mathrm{pa}_{\mathcal{G}_{\mathrm{sum}}}(j)$ denotes the parents of node $j$ in the summary graph, and $a^{j}_{i,t-k}$ is only allowed to be non-zero if $(i,t-k)\to(j,t)$ is an edge in $\mathcal{G}$. SVAR-processes can be considered the time series analogue of additive linear noise models for which sortability was discussed by Reisach et al. (2021; 2023). When generating data from such a model, coefficients are typically randomly drawn, sometimes with a pre-specified proportion of contemporaneous links, and the process is being run until it has converged to a stationary distribution (or is discarded if the distribution is non-stationary).
@@ 2.2 NOTEARS and Derivatives
Zheng et al. (2018) propose the continuous score-based causal discovery method NOTEARS, which embeds the discrete search space of DAGs into a continuous one by using the differentiable function $h(\mathbf{W})=tr\>e^{\mathbf{W}\circ\mathbf{W}}-d$. This function is 0 if and only if $\mathbf{W}$ is the adjacency matrix of an acyclic graph and hence measures the DAGness of $\mathbf{W}$ (Zheng et al., 2018). By combining this function with a score evaluating how well the estimated weight matrix $\mathbf{W}$ fits the data, Zheng et al. (2018) formulate the constrained optimisation problem to find
[…]
@@
The DYNOTEARS algorithm Pamfil et al. (2020) modifies the NOTEARS algorithm to work with time-lagged and auto-correlated dependencies by redefining the optimisation problem to $\min_{\mathbf{W}1}\ell(\mathbf{W_{c}},\mathbf{W_{l}})\text{ s.t. }\mathbf{W}_{c}\text{ is acyclic}$, where $\mathbf{W_{l}}\in\mathbb{R}^{d\times d\times\tau_{max}}$ is the lagged adjacency matrix and $\mathbf{W}_{c}$ is the contemporaneous adjacency of the underlying time series process. As the underlying time series graph to estimate is acyclic if and only if $\mathbf{W}_{c}$ is acyclic, it suffices to enforce the acyclicity constraint only on $\mathbf{W}_{c}$ (Pamfil et al., 2020). To ensure sparsity of $\mathbf{W}$, Pamfil et al. (2020) also add an $\ell_{1}$ penalty term, leading to the constraint optimisation problem
[…]
@@
The continuous optimisation problems of NOTEARS and DYNOTEARS can be solved efficiently by rewriting the problem using the augmented Lagrangian method and using a numerical solver such as L-BFGS (Zheng et al., 2018; Pamfil et al., 2020). After applying the numerical optimisation algorithm in both of the algorithms, a threshold $t$ is applied to remove weights close to zero (Zheng et al., 2018; Pamfil et al., 2020).
@@ 2.3 Sortability Criteria
In the following, we briefly reiterate the two sorting criteria varsortability and $R^{2}$-sortability introduced by Reisach et al. (2021) and Reisach et al. (2023) respectively. In essence, both of these approaches calculate a score $s$ of sortability in a comparable manner; $s$ is the measurement of the degree of agreement between the true causal order and the increasing order of a sortability criterion cricri.
For any causal model containing the variables $\{X^{(1)},\ldots,X^{(d)}\}$ with a (non-degenerate) adjacency matrix $\mathbf{W}$, the sortability score is the fraction of directed paths that start from a node with a strictly lower sortability criterion than the node they end in. Thus the sortability for one selected criterion cricri can be calculated as
@@
The matrix $\mathbf{W}$ in Equation 4 is set to the power $k$, since the $(i,j)$ entry of the $k$-th power of an adjacency matrix of a DAG exactly counts the number of directed paths from $i$ to $j$. In the case of varsortability Reisach et al. (2021), the criterion $cri(X^{i})=\mathrm{Var}(X^{i})$ is the marginal variance, whereas in the case of $R^{2}$-sortability, it is the coefficient of determination $cri(X^{i})=R^{2}(X^{i})$ which acts as a proxy for the fraction of the variance of $X_{i}$ that is explained by its causal parents, (see Reisach et al. (2023) for details). Varsortalility $v$ is defined by using the marginal variance as cricri. $R^{2}$-sortability $r$ is defined by using the obtained $R^{2}$-coefficients as the sorting criterion cricri (Reisach et al., 2021; 2023).
Reisach et al. (2021) showed that varsortability is usually high in data generated by LANMs (see the analytical and empirical proof in their paper). Based on their sortability criteria, Reisach et al. (2021) and Reisach et al. (2023) introduced the baseline methods varsortnregress and $R^{2}$-sortnregress respectively. These algorithms sort the system variables based on their variances or $R^{2}$-scores which are estimated by fitting a regression model; these simple benchmark methods are then shown to have similar performance to some state of the art causal discovery algorithms (Reisach et al., 2021; 2023) on LANM data.
@@ 3 Modified Sortability Criteria for Summary Graphs
However, any pair of processes $X^{i}$ and $X^{j}$ that belong to the same strongly connected component of the summary graph (i.e. that are connected by a cycle) would always contribute a $1$ to the numerator and a $2$ to the denominator of Equation 4.
@@
$s:=\frac{\sum_{(i,j)\in\mathcal{AP}(\mathcal{G}_{\mathrm{sum}})}increasing(cri(X_{i}),cri(X_{j}))}{\sum_{(i,j)\in\mathcal{AP}(\mathcal{G}_{\mathrm{sum}})}1}\in[0,1],$ (6)
@@
CCBBDDAACri(C)=3Cri(C)=3Cri(B)=2Cri(B)=2Cri(D)=2.5Cri(D)=2.5Cri(A)=1Cri(A)=1
Figure 1: Example of the calculation of $s$ with cycles.
[…]
@@
Based on heuristics and empirical findings, Reisach et al. (2021) argue that high variability causes gradient-based optimisation algorithms such as NOTEARS(Zheng et al., 2018) to favour graphs whose edges point in the direction of increasing marginal variances during their first optimisation step. Since DYNOTEARS is a time-series adaptation of NOTEARS, that essentially coincides with NOTEARS on the contemporaneous components of a ts-graph, Reisach et al. (2021)’s arguments might be applicable to (the contemporaneous part of) DYNOTEARS as well.
However, in the recent work Ng et al. (2024) provide examples in which continuous optimization-based methods can not perform well in the presence of high varsortability. Ng et al. (2024) also give an alternative explanation for why continuous structure learning performs significantly worse after standardisation of LANM data. Continuous structure learning assumes equal noise variances for all variables in the system, which is violated in the standardised model leading to poor performance.